# LLM-based Continuous Intrusion Detection Framework for Next-Gen Networks

### v1

@@ Abstract
Cybersecurity, Evolving Threats, Continuous Detection, Transformer Encoder.
@@ I Introduction
Network systems are becoming increasingly complex due to ongoing technological advancements. Next generations of network are expected to enable various interactions and exchanges over the Internet, supporting services such as digital brain-computer interfaces, in-body health networks, and extended reality. This transition will be underpinned by achieving data rates around terabits per second with ultra-low latency. This continuous evolution and expansion of network infrastructure lead to an extremely heterogeneous environment characterized by massive volumes of data. This can complicate network monitoring, exposing systems to malicious activities. Furthermore, as new services emerge from network improvements, sophisticated threats are likely to increase correspondingly, becoming harder to detect [8]. Therefore, it is imperative to develop robust security measures within these highly heterogeneous ecosystems. Various machine learning methods have been explored to enhance security across diverse network environments, including smart grids [16], virtual private networks [13], and IoT ecosystems [14]. Among these methods, Support Vector Machines (SVMs) are widely utilized for their ability to classify data into distinct categories, making them effective for detecting intrusions and anomalies within network traffic. Boosting methods enhance the performance of weak classifiers by combining them into a strong ensemble classifier, thereby improving detection rates for malicious activities. Deep Neural Networks (DNNs) leverage multiple layers of neurons to learn intricate patterns in data, which proves particularly useful for identifying advanced threats in large-scale networks. Recurrent Neural Networks (RNNs) are specifically designed to analyze sequences of network packets, enabling the identification of patterns indicative of cyber threats. Long Short-Term Memory (LSTM) networks excel at detecting temporal patterns in network traffic and are frequently employed in applications such as detecting Distributed Denial of Service (DDoS) attacks [9]. In addition to their diversity, the data in these ecosystems are dynamic, presenting the challenge of encountering unknown data flows during analysis. Machine learning-empowered intrusion detection systems (IDSs) often struggle to keep pace with the constant evolution of network flows, as they are typically trained to detect a fixed or predefined set of attacks [20]. However, in real-world scenarios, intrusion data are collected incrementally. As new data emerge, these models may suffer from catastrophic forgetting, a prevalent issue in machine learning where previously learned classes are forgotten when training on new ones [7]. Recently, techniques involving LLMs have demonstrated an ability to effectively manage vast volumes of data while understanding the underlying context within them. Given the unknown pattern in next generation of networks, leveraging LLMs presents a promising approach for efficiently processing and analyzing such data. Studies have shown that LLMs can address various challenges in telecommunications, such as optimizing the reward process in reinforcement learning [6], and providing zero-shot image classification in complex signal transmissions [22]. However, it is important to note that LLMs were not initially designed for these applications; their primary utility lies in natural language processing (NLP). Consequently, directly applying them to domain-specific tasks can be challenging, necessitating fine-tuning for optimal performance. To the best of our knowledge, this work proposes the first hybrid incremental intrusion detection framework that leverages LLMs to address emerging threats.
@@
  • We investigate prior works related to intrusion detection, focusing on continuous detection.
  • We propose a framework that leverages language models to efficiently capture contextual information within network traffic.
  • We evaluate the continuous detection capability of our framework on a popular IDS dataset, highlighting its effectiveness.
@@ II Related Work
Recent advances in large language models (LLMs) have shown significant potential in efficiently processing natural language. In our prior work [2], we demonstrated that it is feasible to deploy an LLM on resource-constrained devices in a decentralized context while maintaining good overall accuracy.
However, BERT’s static training on a fixed set of attacks introduces uncertainty when new attack flows emerge, as it tends to classify novel attacks based on patterns from previously known categories. Similarly, [8] proposed a proactive network intrusion prediction framework that integrates fine-tuned GPT and BERT models with an LSTM classifier to predict and evaluate IoT network traffic. While effective, this framework primarily addresses the binary nature of network traffic. In an effort to enhance security in 6G architectures, [17] suggested combining zero-trust architecture with artificial intelligence by deploying AI-based agents to protect the network from internal, external, and zero-day attacks. However, their approach lacks a focus on detailed analysis of newly detected threats and does not utilize advanced AI models for in-depth flow analysis—an essential feature as threats become increasingly sophisticated.
@@ III Proposed framework
In this section we present some popular datasets used in IDS and focus on CSE-CIC-IDS2018, we then expose the preprocessing methods we employed, and the design of the model we used as the baseline for our framework. After that we present the method used to ensure incremental learning.
@@ IV-B Model design
LLMs are typically used for tasks such as sentiment analysis, translation, and question answering. Recent studies, however, have begun exploring their application in attack detection. To achieve our objective, we developed the baseline model shown in Figure 2. Specifically, it consists of fine-tuning BERT, using only a few layers with a softmax activation at the output, enabling it to learn contextual representations tailored to attack detection.
In its basic architecture, BERT has two free parameters: $H$, the size of the hidden layers, and $L$, the number of encoder layers in the stack. From these, the size of the intermediate layer and the number of attention heads are derived.
By retaining the first four layers, which primarily capture the syntactic structure, we reduced the model’s size by approximately 90%, making it suitable for our framework [2].
@@ V-C Results and discussion
(a) Performance of the Binary Detector
(b) Initial performance of the contious learning model
Figure 3: Evaluation of the binary detector and the initial continous learning model performance
(a) chunk 1: DoS attacks Hulk and DDoS attacks HOIC
(b) chunk 2: SSH Brute-force, FTP Brute-force, and Infiltration
Figure 4: Silhouette Analysis for Determining the Optimal Number of Clusters in Different Attack Types
(a) t-SNE 2D visualization of chunk 1 embeddings
@@
Figure 5: t-SNE 2D visualization of chunk embeddings for two different data chunks
(a) Confusion matrix after adding clusters from chunk 1.
(b) Confusion matrix after adding clusters from chunk 2.
Figure 6: Granularity in the confusion matrix after incrementally adding newly detected clusters from chunk 1 and chunk 2.
@@
The evaluation process begins with assessing the detection model’s capability to accurately discriminate between malicious and benign traffic. As illustrated in Figure 3(a), the model achieves a perfect classification, evidenced by the confusion matrix showing no false positives or false negatives. This indicates that the detector can reliably identify malicious traffic, achieving a recall of 100%. This performance is critical for ensuring the security of network systems, as it signifies that the model does not misclassify harmful activities as benign, which could lead to significant security breaches.
@@
Next, as shown in Figure 3(b), we assess the identification model using known attack patterns, specifically targeting DDoS attacks-LOIC-HTTP and Bot. During this evaluation, the model effectively identifies these known attacks, demonstrating its robustness. The class ”0_other” plays a crucial role in establishing the decision boundary between known and unknown traffic, allowing the model to distinguish between familiar and unfamiliar patterns.
@@
Subsequently, we introduce two batches of unknown attack traffic to rigorously test the model’s ability to detect new or unseen attacks. Chunk 1 contains DoS attacks-Hulk and DDoS attack-HOIC, while chunk 2 consists of SSH brute-force, FTP brute-force, and infiltration attacks. To manage the complexity posed by these unknown attacks, we leverage the Gaussian Mixture Model (GMM). Given that the data features are represented as BERT embeddings in high-dimensional space, this probabilistic approach efficiently estimates the likelihood of a data point belonging to a specific cluster. This operation is performed iteratively until new unknown traffic is discovered, ensuring the model adapts to evolving threat landscapes.
To determine the optimal number of clusters for separating the unknown attacks in each batch, we utilize the silhouette score, as illustrated in Figure 4. Upon introducing chunk 1, the identification model flags this traffic as unknown. Analysis in Figure 4(a) reveals that the optimal number of clusters is determined to be $4$. This indicates that the traffic comprising DoS attacks-Hulk and DDoS attack-HOIC exhibits enough diversity to be classified into four distinct patterns. Consequently, we extend the classification layer with four new nodes to accommodate these newly identified clusters, and the model is updated accordingly.
For chunk 2, which includes SSH brute-force, FTP brute-force, and infiltration attacks, the silhouette score depicted in Figure 4(b) suggests an optimal number of clusters of $3$. This number corresponds to the distinct behaviors exhibited by the different attack types, reinforcing the model’s adaptability and precision in identifying unique attack patterns. Similar to chunk 1, the classification layer is extended, and the model is updated to incorporate these new behaviors.
Additionally, the visualization of these clusters is presented in Figure 5, where each cluster is distinguished using t-SNE to project high-dimensional data into two dimensions. In Figure 5(b), some blue points appear above the red and yellow points, which does not necessarily indicate incorrect clustering; in three dimensions, these clusters are well-separated.
In Figure 6, we observe the model’s precision after incrementally adding newly discovered clusters. Specifically, in Figure 6(a), a decrease in accuracy is noted for DDoS attacks-HTTP, which shifts in favor of $X1\_1$, the first cluster from chunk 1. This phenomenon is likely attributed to shared features within the feature space, where overlapping characteristics between attack types lead to slight misclassifications. A similar, though less pronounced, effect is observed with the other clusters $X1\_2$, $X1\_3$, and $X1\_4$. Despite these minor accuracy drops, the model maintains robust overall performance, reinforcing its reliability in identifying known threats.
In Figure 6(b), after the integration of clusters from chunk 2, we observe a pattern consistent with the previous results. Notably, there is an improvement in accuracy for DDoS attacks-HTTP, reflecting the model’s adaptive learning capabilities. However, a slight degradation is evident in newly introduced classes such as $X2\_1$ and $X2\_2$, where a small submatrix of misclassifications is noticeable. This observation highlights the inherent challenge in balancing the model’s ability to adapt to new patterns while maintaining accuracy for previously learned classifications.
Figure 7 illustrates the impact of exposing our Intrusion Detection System (IDS) to new attacks on various detection metrics, emphasizing the evolution of these metrics as the system encounters previously unseen threats. The ”baseline” represents the model’s performance on already known attacks. The introduction of chunk 1 is followed by an evaluation of identification and classification, which is then repeated with chunk 2. The baseline demonstrates strong performance across all metrics, indicating that the model is thoroughly familiar with the traffic patterns and can accurately classify known attacks.
@@
Upon introducing chunk 1, we observe a slight drop in performance across all metrics, underscoring the challenges associated with integrating previously unseen attacks. Despite this decrease, the performance remains high due to the effectiveness of the clustering technique, which successfully captures the variations within chunk 1. The model’s performance decreases again with the addition of chunk 2, as anticipated, given that this batch includes more complex and diverse attack types. Nonetheless, the overall performance in detecting both known and newly added attacks remains high, demonstrating the model’s resilience and effectiveness in maintaining security against evolving threats.
@@ VI CONCLUSION
In this article, we proposed an adaptive framework aimed at continuously detecting emerging attacks in networks to enable prompt and efficient responses, thereby preserving the quality of service. The framework leverages a transformer encoder architecture to learn hidden patterns in a bidirectional manner within network traffic, achieving remarkable performance in distinguishing attacks from legitimate traffic.
The second phase of the process focuses on incrementally identifying different types of attacks within the malicious traffic. By utilizing the Gaussian Mixture Model (GMM), the framework successfully clusters features extracted from the unknown traffic, which are then used to update the identification module. Even after adding all discovered clusters, the framework maintains high detection accuracy, with identification accuracy and recall remaining at 95.6%. Notably, the framework retains knowledge of past classes, which enhances its ability to differentiate between known and new threats.
Our ultimate goal is to build a robust and scalable system capable of real-time intrusion detection, leveraging the promising approach of adaptive models to address evolving network threats. Future work will investigate efficient ways to update the model without relying on past data, thereby improving its adaptability and performance in dynamic environments.

### v2

@@ Abstract
Cybersecurity, Evolving Threats, Continuous Detection, language models.
@@ I Introduction
Network systems are becoming increasingly complex due to ongoing technological advancements. Next generations of network are expected to enable various interactions and exchanges over the Internet, supporting services such as digital brain-computer interfaces, in-body health networks, and extended reality. This transition will be underpinned by achieving data rates around terabits per second with ultra-low latency. This continuous evolution and expansion of network infrastructure lead to an extremely heterogeneous environment characterized by massive volumes of data. This can complicate network monitoring, exposing systems to malicious activities. Furthermore, as new services emerge from network improvements, sophisticated threats are likely to increase correspondingly, becoming harder to detect [8]. Therefore, it is imperative to develop robust security measures within these highly heterogeneous ecosystems. Various machine learning methods have been explored to enhance security across diverse network environments, including smart grids [16], virtual private networks [13], and IoT ecosystems [14]. Among these methods, Support Vector Machines (SVMs) are widely utilized for their ability to classify data into distinct categories, making them effective for detecting intrusions and anomalies within network traffic. Boosting methods enhance the performance of weak classifiers by combining them into a strong ensemble classifier, thereby improving detection rates for malicious activities. Deep Neural Networks (DNNs) leverage multiple layers of neurons to learn intricate patterns in data, which proves particularly useful for identifying advanced threats in large-scale networks. Recurrent Neural Networks (RNNs) are specifically designed to analyze sequences of network packets, enabling the identification of patterns indicative of cyber threats. Long Short-Term Memory (LSTM) networks excel at detecting temporal patterns in network traffic and are frequently employed in applications such as detecting Distributed Denial of Service (DDoS) attacks [9]. In addition to their diversity, the data in these ecosystems are dynamic, presenting the challenge of encountering unknown data flows during analysis. Machine learning-empowered intrusion detection systems (IDSs) often struggle to keep pace with the constant evolution of network flows, as they are typically trained to detect a fixed or predefined set of attacks [20]. However, in real-world scenarios, intrusion data are collected incrementally. As new data emerge, these models may suffer from catastrophic forgetting, a prevalent issue in machine learning where previously learned classes are forgotten when training on new ones [7]. Recently, techniques involving LLMs have demonstrated an ability to effectively manage vast volumes of data while understanding the underlying context within them. Given the unknown pattern in next generation of networks, leveraging LLMs presents a promising approach for efficiently processing and analyzing such data. Studies have shown that LLMs can address various challenges in telecommunications, such as optimizing the reward process in reinforcement learning [6], and providing zero-shot image classification in complex signal transmissions [22]. However, it is important to note that LLMs were not initially designed for these applications; their primary utility lies in natural language processing (NLP). Consequently, directly applying them to domain-specific tasks can be challenging, necessitating fine-tuning for optimal performance. To the best of our knowledge, this work proposes the first hybrid incremental intrusionTo the best our knowledge, this work proposes the first hybrid continuous intrusion detection framework that utilizes LLMs to address emerging threats. The code of our simulation is available  [^1] https://github.com/freddyAdjh/LLM-CIDF-next_gen.
@@
  • We present a novel framework that leverages language models to detect and identify malicious traffic, effectively capturing contextual information with high accuracy.
  • We propose an adaptive clustering approach based on a Gaussian Mixture Model applied to network flow embeddings, enabling the dynamic identification and categorization of previously unknown attack types for continuous learning and model updating.
  • We demonstrate the framework’s effectiveness in adapting to emerging threat patterns, achieving high accuracy and recall in detection and identification.
@@ II Related Work
Recent advances in LLMs have shown significant potential in efficiently processing natural language. In our prior work [2], we demonstrated that it is feasible to deploy an LLM on resource-constrained devices in a decentralized context while maintaining good overall accuracy.
However, BERT’s static training on a fixed set of attacks introduces uncertainty when new attack flows emerge, as it tends to classify novel attacks based on patterns from previously known categories. Similarly, [8] proposed a proactive network intrusion prediction framework that integrates fine-tuned GPT and BERT models with an LSTM classifier to predict and evaluate IoT network traffic. While effective, this framework primarily addresses the binary nature of network traffic. In an effort to enhance security in 6G architectures, [17] suggested combining zero-trust architecture with artificial intelligence by deploying AI-based agents to protect the network from internal, external, and zero-day attacks. However, their approach lacks a focus on detailed analysis of newly detected threats and does not utilize advanced AI models for in-depth flow analysis, an essential feature as threats become increasingly sophisticated.
@@ III Proposed framework
In this section we present some popular datasets used in IDS and focus on CSE-CIC-IDS2018, we then expose the preprocessing methods we employed, and the design of the model we used as the baseline for our framework.
@@ IV-B Model design
LLMs are commonly used for tasks such as sentiment analysis, translation, and question answering. However, recent research has demonstrated their potential in detecting network attacks. To leverage this capability, we fine-tuned a BERT model, focusing on its initial layers to efficiently capture the syntactic structure of network flow data.
Specifically, as illustrated in Figure 2, we retained the first four layers of BERT approximately 10% of the original model which are primarily responsible for understanding the grammar and syntax of the input data. This approach significantly reduced the model’s complexity and computational cost, making it suitable for our purposes.
After processing the network flow data through these layers, a softmax function was applied to generate a probability distribution over different attack classes. This enables the model to classify network flows into specific attack categories with high confidence.
@@ V-C Results and discussion
Figure 3: Performance of the Binary Detector
Figure 4: Initial performance of the contious learning model
(a) Silhouette score for optimal cluster selection of Chunk 1
(b) t-SNE 2D visualization of Chunk 1 embeddings
(c) Confusion matrix after adding clusters from Chunk 1
Figure 5: Evaluation of clustering results for Chunk 1: silhouette score, t-SNE visualization, and confusion matrix
(a) Silhouette score for optimal cluster selection of Chunk 2
@@
(c) Confusion matrix after adding clusters from Chunk 2 to the updated model
Figure 6: Evaluation of clustering results for Chunk 2 on the updated model: silhouette score, t-SNE visualization, and confusion matrix
@@
The evaluation process starts with the binary detection model, as illustrated in Figure 3, it achieves a perfect classification, evidenced by the confusion matrix showing no false positives or false negatives. This indicates that the detector can reliably identify malicious traffic, achieving a recall of 100%. This performance is critical for ensuring the security of network systems, as it signifies that the model does not misclassify harmful activities as benign, which could lead to significant security breaches.
@@
Next, as shown in Figure 4, we assess the identification model using known attack patterns, specifically targeting DDoS attacks-LOIC-HTTP and Bot. During this evaluation, the model effectively identifies these known attacks, demonstrating its robustness. The class ”0_other” plays a crucial role in establishing the decision boundary between known and unknown traffic, allowing the model to distinguish between familiar and unfamiliar patterns.
@@
Subsequently, we introduce two batches of unknown attack traffic to rigorously test the model’s ability to detect new or unseen attacks. Chunk 1 contains DoS attacks-Hulk and DDoS attack-HOIC, while chunk 2 consists of SSH brute-force, FTP brute-force, and infiltration attacks. To manage the complexity posed by these unknown attacks, we leverage the Gaussian Mixture Model. Given that the data features are represented as BERT embeddings in high-dimensional space, this probabilistic approach efficiently estimates the likelihood of a data point belonging to a specific cluster. This operation is performed once until new unknown traffic is discovered, ensuring the model adapts to evolving threat landscapes.
The following results provide insights into the model’s ability to continuously learn and accurately identify and classify attack patterns over time. In Figure 5(a), the silhouette score for chunk 1 indicates that the optimal number of clusters is 4. This suggests that the traffic consisting of DoS attacks-Hulk and DDoS attacks-HOIC exhibits sufficient diversity to be grouped into four distinct patterns. After clustering, the classification layer is extended with four additional nodes to accommodate the new clusters, enabling the model to update its structure accordingly. This step is crucial for maintaining the model’s flexibility. The t-SNE 2D visualization in Figure 5(b) further supports this clustering result by projecting the high-dimensional data into two dimensions. As shown, the clusters are well-separated in the 2D space, confirming the effectiveness of the chosen number of clusters. Figure 5(c) illustrates the impact of adding the newly discovered clusters from chunk 1 to the updated baseline model. A decrease in accuracy is observed, particularly for DDoS attacks-HTTP, which shifts towards the first cluster $X1\_1$. This can be attributed to the overlapping features within the feature space, where shared characteristics between attack types lead to minor misclassifications. Nevertheless, the classification model continues to perform robustly, retaining high accuracy for most attack patterns.
In contrast, the clustering results for chunk 2, shown in Figure 6(a), suggest that the optimal number of clusters is 3. This corresponds to the distinct behaviors of the various attack types in this chunk. As with chunk 1, the classification layer is again extended after clustering to update the model with the newly identified clusters. The t-SNE visualization in Figure 6(b) presents the clusters from chunk 2, where the separation is clear. Some blue points, which represent a particular attack type, appear to overlap with the red and yellow points in the 2D plot, but this does not imply misclassification. Rather, this overlap is resolved when viewed in three dimensions, where these clusters are better separated, reinforcing the robustness of the clustering method. Figure 6(c) shows the model’s performance after incorporating the clusters from chunk 2. We observe an improvement in accuracy for DDoS-HTTP attacks, indicating the model’s enhanced ability to differentiate between attack types after retraining. This improvement highlights the model’s capacity to adapt to new data, as retraining allows it to update its decision boundaries and refine its classifications based on recent information. However, a slight accuracy degradation is noted for the newly learned clusters, such as $X2\_1$ and $X2\_2$, where a small misclassification submatrix appears. This highlights the challenge of maintaining the same level of precision when new attacks are introduced.
@@
To assess the model’s ability to evolve over time in the face of unknown attacks, we use the previously mentioned metrics applied to three simulated use cases: (i) The baseline, where the IDS is familiar with all attack types; (ii) the baseline with chunk 1, where the IDS encounters unknown attacks and uses clustering to segment latent unknown attack types, and updates the model, and (iii) the baseline with chunk 1 and chunk 2, where the model has already been updated with attacks from chunk 1 and faces new attacks in chunk 2. In each case, identification, classification, and model updates follow the same steps as for chunk 1. Figure 7 illustrates the impact of exposing the Intrusion Detection System (IDS) to new attacks across various detection metrics, emphasizing the evolution of these metrics as the system encounters previously unseen threats. The ”baseline” represents the model’s performance on known attacks. The introduction of chunk 1 is followed by an evaluation of identification and classification, which is then repeated with chunk 2. The baseline demonstrates strong performance across all metrics, indicating that the model is well-equipped to classify known attacks. Upon introducing chunk 1, a slight drop in performance is observed across all metrics, highlighting the challenges of integrating previously unseen attacks. Despite this decrease, the performance remains high due to the effectiveness of the clustering technique, which successfully captures the variations within chunk 1. A further decrease in performance is observed with the addition of chunk 2, as expected, given that this batch includes more complex and diverse attack types. Nonetheless, the overall performance in detecting both known and newly added attacks remains high, showcasing the model’s resilience and its ability to maintain security against evolving threats.
@@ VI CONCLUSION
In this article, we propose an adaptive framework designed to detect, identify and classify emerging next-generation network attacks. The framework employs a transformer encoder architecture to capture intricate bidirectional patterns in network traffic, achieving strong performance in distinguishing attacks from legitimate activity. Once malicious traffic is detected, the focus shifts to the continuous identification of diverse attack types within this traffic. Leveraging the Gaussian Mixture Model (GMM), the framework effectively clusters features from unknown traffic, enabling dynamic updates to the identification module. Even after integrating new clusters, the framework sustains high detection accuracy, with classification accuracy and recall rates maintaining at 95.6%. Notably, the model preserves knowledge of previously learned classes, enhancing its capability to differentiate known threats from new ones.
Future research will explore efficient model update techniques that do not rely on prior data, further improving adaptability and performance in dynamic environments.